General Announcement:
- A. Galler informs the beam delivery start is delayed currently up to 15:00 due to difficulties in setting up the machine after the FD week.
- Couple of meetings regarding the LIMP are scheduled; plans have been distributed and will be shared during the meetings.
Please attend the meetings and if you want an invitation, contact Martin Doering. Also a DATA meeting on LIMP is happening today and please contact Patrick G., Steffen H. or Martin Doering for invitation.
CTRL:
- Nothing to report.
DA:
- Not present.
EEE:
- Nothing to report.
VAC:
- Nothing to report.
XPD:
- Nothing to report; all fine.
DET:
- Nothing major to report.
ITDM:
- Nothing to report.
XRO:
- Not present.
LAS:
- Nothing to report.
UNSYS:
- Nothing to report; everything fine.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Nothing to report.
HED:
- Nothing to report.
SCS:
- Preparing for next weeks.
SQS:
- Nothing to report.
SXP:
- Nothing to report. |